OpenStudy (anonymous):

Darlene kicks a soccer ball off the ground and in the air with an initial velocity of 34 feet per second. Using the formula H(t) = −16t^2 + vt + s, what is the maximum height the soccer ball reaches? 17.7 feet 18.1 feet 19.3 feet 20.2 feet

OpenStudy (anonymous):

i will use calculus based approach

OpenStudy (mimi_x3):

well there is a calculus method; if you know calculus.. differentiate it

OpenStudy (anonymous):

Calculate maximum of this function

OpenStudy (anonymous):

take derivative to get critical point

OpenStudy (anonymous):

derivative is zero on maximum

OpenStudy (anonymous):

h'(t)=-32t+v

OpenStudy (anonymous):

set equal to zero -32t+v=0

OpenStudy (anonymous):

but v=34 ft/sec given -32t+34=0

OpenStudy (anonymous):

32t=34

OpenStudy (anonymous):

soo.,.

OpenStudy (anonymous):

t=34/32=1.0625

OpenStudy (anonymous):

put t=1.0625 in the original equation

OpenStudy (mimi_x3):

i think there is a typo in the formula..

OpenStudy (anonymous):

@Mimi_x3 let me know i do a lot typo mistakes :)

OpenStudy (anonymous):

@rebecca1233 can you do this now ?? just put value of t

OpenStudy (anonymous):

can you help me so i put 1.0625 in the t spot

OpenStudy (mimi_x3):

not you; the asker \(h(t) = -16t^2 +vt +s\) there are two unknown variables..

OpenStudy (anonymous):

−16t^2 + vt + s,

OpenStudy (anonymous):

ok

OpenStudy (mimi_x3):

and if you sub back the t the \(s\) is unknown

OpenStudy (anonymous):

ugh can you guys help me with this ..

OpenStudy (anonymous):

h(t)=-16(1.0625)^2+34(1.0625+0=

OpenStudy (mimi_x3):

then \(s\) is a constant? and you assume its \(0\)?

OpenStudy (anonymous):

let me know the answer @rebecca1233 just calculate the above.

OpenStudy (anonymous):

ok i did it =18.1 feet . use your calculator.

OpenStudy (anonymous):

@rebecca1233